OpenStudy (anonymous):

z1=-3+4i z2=2-i a+bi form z1/z2=?

OpenStudy (anonymous):

did you find out what a conjugate is yet?

OpenStudy (anonymous):

Changing the sign in the middle of two terms?

OpenStudy (mertsj):

Why did you post this again? I showed you how to do it.

OpenStudy (anonymous):

Yes, but my options didn't match up. It's a multiple choice question.

OpenStudy (mertsj):

Why did you get when you multiplied by the fraction I showed you?

OpenStudy (anonymous):

-15... Something over something * something :/ I forget.

OpenStudy (mertsj):

\[\frac{-3+4i}{2-i}\times\frac{2+i}{2+i}=\]

OpenStudy (mertsj):

Multiply the numerators. What do you get?

OpenStudy (anonymous):

I think I got it. -2+i

OpenStudy (anonymous):

Thanks for you help!

OpenStudy (mertsj):

yw